# QMetro++ - Python optimization package for large scale quantum metrology with customized strategy structures

### v1

@@ 2.4 Iterative see-saw
In the ISS optimization method developed in the quantum metrology context in [40, 41, 42, 43, 17, 44, 22] one considers the pre-QFI function:
@@
where $\rho_{\theta}=\Lambda_{\theta}\otimes\mathbb{I}_{\mathcal{A}}(\rho_{0})$, and where the dimension of $\mathcal{A}$ can be chosen arbitrarily and this choice may affect the final result, unlike in the MOP method. It turns out that for a given $\rho_{0}$ the maximization of $F$ over hermitian matrices $\mathfrak{L}$ yields QFI of the state $\rho_{\theta}$:
@@
with the solution $\mathfrak{L}^{\diamond}$ being equal to SLD matrix $L$ of $\rho_{\theta}$ [41]—hence we will use the name pre-SLD matrix for $\mathfrak{L}$. It follows that the channel QFI (9) is simply a maximization of $F$ over both of its arguments:
@@
This problem can be solved by initializing $\rho_{0}$ and $\mathfrak{L}$ at random and then by interchangeably making one argument constant and maximizing the other one. Each such optimization will increase the value of $F$ and we can proceed until the change of $F$ in a number of consecutive iterations becomes smaller then some established precision $\epsilon$. In case of the single channel optimization the procedure is implemented in iss_channel_qfi.
ISS algorithm can be easily applied to parallel strategy by simply substituting $\Lambda_{\theta}$ with $\Lambda^{(N)}_{\theta}=\Lambda^{\otimes N}_{\theta}$—this is implemented in iss_parallel_qfi function. This approach, however, suffers from the same problem as MOP—namely, it has an exponential complexity in $N$.
@@
|$\ket{\psi}=\sum_{a_{0},...,a_{N}}\psi^{a_{0}...a_{n}}\ket{a_{0}...a_{N}}\in\mathcal{I}^{(N)},$ (30)
can be expressed as a tensor network called matrix product state (MPS) [45, 46]:
@@
where indices $a_{i}$ are called physical and horizontal indices that connect tensors $\psi_{i}$ are called bond. Range of the bond indices called bond dimension and denoted by $r_{\mathrm{MPS}}$ controls the strength of possible correlations between sites of |$\ket{\psi}$, $r_{\mathrm{MPS}}=1$ means that the state is separable and $r_{\mathrm{MPS}}\approx\dim\mathcal{I}^{(N)}$ allows for arbitrary correlations. Analogously, pre-SLD matrix can be expressed as a matrix product operator (MPO):
@@
with bond dimension $r_{\mathfrak{L}}$. We can then write pre-QFI function as
@@
Just like in the case of a single channel optimization, we can maximize the above function over one argument at a time while others stay constant. Then after optimization of each argument, we repeat the procedure until the value of $F$ converges. Note that if $r_{\mathfrak{L}}$ is too small, then the optimization over pre-SLD will not converge to the SLD matrix and value of pre-QFI will not be equal to QFI.
Nevertheless, the obtained value of $F$ can be always interpreted as a CFI for a measurement given by eigenvectors of $\mathfrak{L}$ [41]. This fact might be used to estimate the maximal value of CFI for local measurement, i.e. measurement in separable basis, because values of pre-QFI for $r_{\mathfrak{L}}=1$ and $r_{\mathfrak{L}}=2$ bound the maximal local measurements CFI from below and above respectively. Procedure employing MPS and MPO to the optimization of parallel strategy is implemented in iss_tnet_parallel_qfi.
For adaptive strategy we can apply the formula on $\rho_{\theta}$ from Fig. 3 to (27) and redefine pre-QFI function as
@@
Then, the QFI for adaptive strategy becomes a maximization of (35) over hermitian $\mathfrak{L}$ and quantum combs $C$ (19)—this is implemented in iss_adaptive_qfi. Once more, this problem can be solved more efficiently if we split arguments of $F$ into smaller parts. This time, however, tensor network of C has a natural interpretation as an input state $\rho_{0}$ and a series of control operations $\mathrm{C}_{i}$ like in Fig. 1(C). Thus we can optimize
@@
We implemented this procedure in iss_adaptive_tnet_qfi.
@@
Let us conclude this section with comparison of MOP and ISS. MOP is a deterministic algorithm with optimality of the numerical solution guaranteed by formal proofs. ISS initializes the parameters of pre-QFI function at random, thus it is nondeterministic. Additionally, there is no guarantee that it converges to the true optimum of the pre-QFI function. The optimality can be validated in some cases by comparing its result with MOP or upper bounds, see Sec. 2.5. It might be also useful to repeat the computation several times and check whether the returned values vary.
On a positive side, what is guaranteed in ISS is that the figure of merit will never decrease in every individual iteration step and that the final result will always be a correct lower bound of a truly optimal QFI.
Moreover, ISS is much more efficient than MOP. Its exact time complexity depends on the strategy and it is hard to estimate. Our numerical data shows that for strategies that are similar to parallel or adaptive strategy and with the use of tensor networks the time complexity is quadratic in $N$. This is significantly better than exponential complexity of MOP and allows to compute QFI for much bigger values of $N$ ($N\leq 5$ for MOP vs $N\leq 100$ for ISS). Moreover, in ISS we can control additional parameters like ancilla or bond dimension and consider strategies different from parallel and adaptive. Finally, while both methods can be used to get the optimal adaptive or parallel protocols, it is much more straightforward for ISS, as it requires only inspecting the final values of the pre-QFI function optimized arguments.
@@ 2.5 Upper bounds
In Sec. 2.3 we explained that the QFI can be computed through minimization of the norm of $\alpha$ (15) over all Kraus representations $\{K_{\theta,k}\}$ of the channel $\Lambda_{\theta}$. It follows that in order to compute QFI for parallel or adaptive strategy we need to optimize over Kraus representations $\{K^{(N)}_{\theta,k}\}$ of the whole channel $\Lambda^{(N)}_{\theta}=\Lambda^{\otimes N}_{\theta}$ which is exponentially hard in $N$. However, if we only want to obtain the upper-bounds, we may dramatically reduce the complexity of the optimization, and reformulate computation of the bound in a way that only minimization over Kraus representation of the single channel $\Lambda_{\theta}$ is required [12, 21, 28, 13, 31].
In case of parallel strategies this leads to an upper-bound of the form [11, 12, 31]:
@@
This bound allows to immediately exclude the possibility of the Heisenberg scaling in the model—it is enough to show that there exist $h$ for which $\beta(h)=0$, which may be formulated as a simple algebraic condition [12, 18].
In case of adaptive strategy, one may again employ the idea of MOP to bound the maximal increase of QFI when the adaptive strategy is extended by one additional step including one more sensing channel [21]:
@@
Both of these bounds can be computed efficiently irrespectively of the value of $N$ [21]. They are implemented in par_bounds and ad_bounds functions respectively.
Importantly, the bounds for parallel and adaptive strategy are asymptotically (for $N\rightarrow\infty$) equivalent and saturable [18, 21]—there are strategies which asymptotically achieve the optimal value of QFI predicted by the bounds.
@@
In the package, asymptotic bound can be computed using asym_scaling_qfi, a function that automatically determines also the character of the scaling.
Finally, there are methods to compute the bounds also in the case of correlated noise models that have been developed recently in [28]. We do not provide exact formulation of the optimization problem that one needs to solve to find the relevant bound as it is much more involved than in case of uncorrelated noise models discussed above. In the package, the correlated noise bounds may be computed using ad_bounds_correlated in the regime of finite channel uses, and ad_asym_bounds_correlated in order to obtain the asymptotic behavior. Unlike in the uncorrelated noise models, there is no guarantee here that the bounds are saturable. Still, they may be systematically tightened at the expense of increasing the numerical complexity, see Sec. 3.5 for more discussion.
@@ 3 Basic package usage—optimization of standard strategies
In this section we present the basic ways to use the package, that utilize high-level functions from protocols and bounds sub-packages. These functions allow the user to compute and bound the QFI within three standard scenarios: single channel use, parallel and adaptive, by simply specifying the parameter encoding channel and the number of uses.
In the following subsections we will apply these methods to study the problem of phase estimation in presence of two paradigmatic decoherence models: dephasing (d) or amplitude damping (a).
The parameter encoding channels $\Lambda_{\theta}$ that we will consider involve is given by Kraus operators of the form
@@
where $\sigma_{z}$ is a Pauli $z$-matrix, $U_{\theta}$ represents unitary parameter encoding, whereas $K_{k}$ are Kraus operators corresponding to one of the two decoherence models. For simplicity, we will always assume that estimation is performed around parameter value $\theta\approx 0$, so after taking the derivatives of Kraus operators we set $\theta=0$—this does not affect the values of the QFI as finite unitary rotation could always be incorporated in control operations, measurements or state preparation stage and hence will not affect the optimized value of QFI.
@@
[…]
where $p=1$ corresponds to no dephasing, $p=1/2$ to the the strongest dephasing case, where all equatorial qubit states are mapped to the maximally mixed states, while $p=0$ represents a phase flip channel. Equivalently, we may use a different Kraus representation where the two Kraus operators are given by:
@@
which may be interpreted as random rotations by angles ±$\pm\epsilon$ with probability $1/2$—in order for the two representations to match we need to fix $p=\cos^{2}(\epsilon/2)$.
@@
where $p=1$ represents no decoherence case, while $p=0$ full relaxation case, where all the states of the qubit are mapped to the ground |$\ket{0}$ state.
Alternatively, dephasing and amplitude damping models can be considered in a continuous time regime. In this case, the state undergoes a process for some time $t$ which imprints on it an information about the parameter $\omega$ related with $\theta$ by $\theta=\omega t$. The derivative of the state over the time is defined using the Lindbladian operator $\mathcal{L}_{\omega}$:
@@
where $H_{\omega}=\hbar\omega\sigma_{z}/2$ is a Hamiltonian generating the evolution $U_{\theta}$ and $\mathcal{D}$ is a dissipation term. The form of the dissipation term depends on the noise model:
$\mathcal{D}^{(d)}[\rho]=\frac{\gamma^{(d)}}{2}(\sigma_{z}\rho\sigma_{z}-\rho),$ (47)
$\mathcal{D}^{(a)}[\rho]=\frac{\gamma^{(a)}}{2}\left(\sigma_{-}\rho\sigma_{-}^{\dagger}-\frac{1}{2}\{\sigma_{-}^{\dagger}\sigma_{-},\rho\}\right),$ (48)
[…]
$2p-1=e^{-\gamma^{(d)}t},\quad p=e^{-\gamma^{(a)}t}.$ (49)
3.1 Specifying the parameter dependent channel
@@
  • from list of Kraus operators and their derivatives over $\theta$,
  • from CJ matrix and its derivative over $\theta$,
@@
  • from a time-independent Lindbladian, its derivative over $\omega$ and specified evolution time.
@@
11 sqrt(p) * np.identity(2),
12 sqrt(1-p) * sz
@@
40
41# Lindbladian and its derivative.
42omega = 0.0 # Parameter.
43t = 1 # Time.
44# Dephasing strength:
45gamma = -np.log(2*p-1)/t
46def lindblad(rho):
@@
55
56# ParamChannel instance created from
57# Lindbladian:
58channel4 = ParamChannel(
59 lindblad=lindblad,
60 dlinblad=dlindblad,
61 time=t, input_dim=2
62)
Figure 4: (A) Diagram of strategy with single parametrized channel $\Lambda_{\theta}$ and ancilla $\mathcal{A}$. Values of QFI for (B) dephasing (C) amplitude damping, various values of $p$ and different methods: MOP - black $\times$, ISS with $d_{\mathcal{A}}=1$ blue dashed line, ISS with $d_{\mathcal{A}}=2$ red dotted line.
Objects of ParamChannel class can be used to compute $\rho_{\theta}=\Lambda_{\theta}(\rho)$ and $\dot{\rho}_{\theta}=\frac{d}{d\theta}\Lambda_{\theta}(\rho)$, e.g.:
@@
or to obtain CJ operator from Kraus operators and vice versa, e.g.:
@@
3.2 Single channel QFI optimization
Having defined the parameter dependent channel using the ParamChannel class, we may now easily compute the single-channel QFI (9), taking into account that we may also make use of the ancilla, see Fig. 4(A). The channel QFI can be computed using one of the two functions:
  • iss_channel_qfi for ISS method,
  • mop_channel_qfi for MOP method.
Function iss_channel_qfi takes two arguments: channel to compute the QFI for and the dimension of the ancillary system. It returns five items: optimized QFI, list of pre-QFI values per algorithm iteration, density matrix of the optimal input state, SLD matrix and the information whether the optimization converged successfully.

### v2

@@ 2.4 Iterative see-saw
In the ISS optimization method developed in the context of quantum metrology [18, 37, 36, 47, 11, 35, 31] one considers the pre-QFI function:
@@
where $\rho_{\theta}=\Lambda_{\theta}\otimes\mathbb{I}_{\mathcal{A}}(\rho_{0})$, and the dimension of $\mathcal{A}$ can be chosen arbitrarily and this choice may affect the final result, unlike in the MOP method. It turns out that for a given $\rho_{0}$ the maximization of $F$ over Hermitian matrices $\mathfrak{L}$ yields the QFI of the state $\rho_{\theta}$:
@@
with the solution $\mathfrak{L}^{\diamond}$ equal to the SLD matrix $L$ of $\rho_{\theta}$ [37]—hence we will refer to $\mathfrak{L}$ as the pre-SLD matrix for $\mathfrak{L}$. It follows that the channel QFI (9) is simply a maximization of $F$ over both of its arguments:
@@
This problem can be solved by initializing $\rho_{0}$ and $\mathfrak{L}$ at random and then by alternately making one argument constant and maximizing the other one. Each such optimization will increase the value of $F$ and we can proceed until the change of $F$ in a number of consecutive iterations becomes smaller than some established precision $\epsilon$. In case of the single-channel optimization the procedure is implemented in iss_channel_qfi.
ISS algorithm can be easily applied to the parallel strategy by simply substituting $\Lambda_{\theta}$ with $\Lambda^{(N)}_{\theta}=\Lambda^{\otimes N}_{\theta}$—this is implemented in iss_parallel_qfi function. This approach, however, suffers from the same problem as MOP—namely, it has an exponential complexity in $N$.
@@
can be expressed as a tensor network called a matrix product state (MPS) [8, 44]:
@@
[…]
with bond dimension $r_{\mathfrak{L}}$. We can then write the pre-QFI function as
@@
Just like in the case of a single channel optimization, we can maximize the above function over one argument at a time while others stay constant. Then after optimization of each argument, we repeat the procedure until the value of $F$ converges. Note that if $r_{\mathfrak{L}}$ is too small, then the optimization over pre-SLD will not converge to the SLD matrix and value of pre-QFI will not be equal to QFI. Nevertheless, the obtained value of $F$ can always be interpreted as a CFI for a measurement given by eigenvectors of $\mathfrak{L}$ [37]. The procedure employing MPS and MPO to the optimization of parallel strategy is implemented in iss_tnet_parallel_qfi.
For the adaptive strategy we can apply the formula on $\rho_{\theta}$ from Fig. 3 to (27) and redefine pre-QFI function as
@@
Then, the QFI for adaptive strategy becomes a maximization of (35) over Hermitian $\mathfrak{L}$ and quantum combs $C$ (19)—this is implemented in iss_adaptive_qfi. Once more, this problem can be solved more efficiently if we split arguments of $F$ into smaller parts. This time, however, the tensor network of C has a natural interpretation as an input state $\rho_{0}$ and a series of control operations $\mathrm{C}_{i}$ like in Fig. 1(C). Thus we can optimize
@@
We implemented this procedure in iss_tnet_adaptive_qfi.
@@
Let us conclude this section with a comparison of MOP and ISS. MOP is a deterministic algorithm with optimality of the numerical solution guaranteed by formal proofs. ISS initializes the parameters of pre-QFI function at random, thus it is nondeterministic. Additionally, there is no guarantee that it converges to the true optimum of the pre-QFI function. The optimality can be validated in some cases by comparing its result with MOP or upper bounds, see Sec. 2.5. It might be also useful to repeat the computation several times and check whether the returned values vary.
On a positive side, what is guaranteed in ISS is that the figure of merit will not decrease at any individual iteration step and that the final result will always be a correct lower bound of a truly optimal QFI.
Moreover, ISS is much more efficient than MOP. Its exact time complexity depends on the strategy and it is hard to estimate. Our numerical data shows that for strategies that are similar to parallel or adaptive strategy and with the use of tensor networks the time complexity is quadratic in $N$. This is significantly better than exponential complexity of MOP and allows to compute QFI for significantly larger values of $N$ ($N\leq 5$ for MOP vs $N\leq 100$ for ISS). Moreover, in ISS we can control additional parameters like ancilla or bond dimension and consider strategies different from parallel and adaptive. Finally, while both methods can be used to get the optimal adaptive or parallel protocols, it is much more straightforward for ISS, as it requires only inspecting the final values of the pre-QFI function optimized arguments.
@@ 2.5 Upper bounds
In Sec. 2.3 we explained that the QFI can be computed via minimization of the norm of $\alpha$ (15) over all Kraus representations $\{K_{\theta,k}\}$ of the channel $\Lambda_{\theta}$. It follows that in order to compute QFI for parallel or adaptive strategy we need to optimize over Kraus representations $\{K^{(N)}_{\theta,k}\}$ of the whole channel $\Lambda^{(N)}_{\theta}=\Lambda^{\otimes N}_{\theta}$—a task that is exponentially hard in $N$. However, if we only want to obtain the upper bounds, we may dramatically reduce the complexity of the optimization, and reformulate the computation of the bound in a way that only minimization over Kraus representations of the single channel $\Lambda_{\theta}$ is required [16, 32, 30, 17, 29].
In case of parallel strategies this leads to an upper bound of the form [20, 16, 29]:
@@
This bound allows immediate exclusion of the possibility of the Heisenberg scaling in the model—it is enough to show that there exist $h$ for which $\beta(h)=0$, which may be formulated as a simple algebraic condition [16, 50].
In case of the adaptive strategy, one may again employ the idea of MOP to bound the maximal increase of QFI when the adaptive strategy is extended by one additional step including one more sensing channel [32]:
@@
Both of these bounds can be computed efficiently regardless of the value of $N$ [32]. They are implemented in the par_bounds and ad_bounds functions respectively.
Importantly, the bounds for parallel and adaptive strategies are asymptotically (for $N\rightarrow\infty$) equivalent and saturable [50, 32]—there are strategies that asymptotically achieve the optimal value of QFI predicted by the bounds.
@@
In the package, the asymptotic bound can be computed using asym_scaling_qfi, a function that automatically also determines the character of the scaling.
Finally, there are methods to compute the bounds also in the case of correlated noise models that have been developed recently in [30]. We do not provide the exact formulation of the optimization problem that one needs to solve to find the relevant bound as it is much more involved than in case of uncorrelated noise models discussed above. In the package, the correlated noise bounds may be computed using ad_bounds_correlated a finite number of channel uses, and ad_asym_bound_correlated in order to obtain the asymptotic behavior. Unlike in the uncorrelated noise models, there is no guarantee here that the bounds are saturable. Still, they may be systematically tightened at the expense of increasing the numerical complexity, see Sec. 3.5 for a more detailed discussion.
@@ 3 Basic package usage—optimization of standard strategies
In this section we present the basic ways to use the package that utilize high-level functions from protocols and bounds subpackages. These functions allow the user to compute and bound the QFI within three standard scenarios: single-channel, parallel and adaptive strategies, by simply specifying the parameter-encoding channel and the number of uses.
In the following subsections we will apply these methods to study the problem of phase estimation in the presence of two paradigmatic decoherence models: dephasing (d) or amplitude damping (a).
The parameter-encoding channels $\Lambda_{\theta}$ that we will consider are given by Kraus operators of the form
@@
where $\sigma_{z}$ is a Pauli $z$-matrix, $U_{\theta}$ represents unitary parameter encoding, whereas $K_{k}$ are Kraus operators corresponding to one of the two decoherence models. For simplicity, we will always assume that estimation is performed around the parameter value $\theta\approx 0$, so after differentiating the Kraus operators we set $\theta=0$—this does not affect the values of the QFI as finite unitary rotations could always be incorporated in control operations, measurements or the state preparation stage and hence will not affect the optimized value of QFI.
@@
$K^{(d)}_{0}=\sqrt{p}\text{1},\quad K^{(d)}_{1}=\sqrt{1-p}\sigma_{z},$ (43)
where $p=1$ corresponds to no dephasing, $p=1/2$ to the strongest dephasing case, where all equatorial qubit states are mapped to the maximally mixed state, while $p=0$ represents a phase flip channel. Equivalently, we may use a different Kraus representation where the two Kraus operators are given by:
@@
which may be interpreted as random rotations by angles $\pm\epsilon$ with probability $1/2$ each—in order for the two representations to match we need to fix $p=\cos^{2}(\epsilon/2)$.
@@
[…]
Alternatively, dephasing and amplitude damping models can be considered in a continuous time regime. In this case, the state undergoes a process for some time $t$ which imprints on it information about the parameter $\omega$ related to $\theta$ by $\theta=\omega t$. The derivative of the state with respect to time is defined using the Lindbladian operator $\mathcal{L}_{\omega}$:
@@
where $H_{\omega}=\hbar\omega\sigma_{z}/2$ is a Hamiltonian generating the evolution $U_{\theta}$ and $\mathcal{D}$ is a dissipation term. The dissipation term has a general form:
$\mathcal{D}[\rho]=\sum_{i}\gamma_{i}\left(L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}\right),$ (47)
where $\{\cdot,\cdot\}$ is the anticommutator, $L_{i}$ are jump operators that determine the type of dissipation and $\gamma_{i}\geq 0$ are damping rates. In case of dephasing and amplitude damping only one pair $L_{i},\gamma_{i}$ is required and it is respectively:
$L^{(d)}=\sigma_{z}/\sqrt{2},$ $2p-1=e^{-\gamma^{(d)}t},$ (48)
$|,\displaystyle L^{(a)}=\ket{0}\bra{1}$, $p=e^{-\gamma^{(a)}t}.$ (49)
3.1 Specifying the parameter-dependent channel
@@
  • from a list of Kraus operators and their derivatives with respect to $\theta$,
  • from a CJ matrix and its derivative over $\theta$,
@@
  • from a Lindbladian, its derivative with respect to $\omega$ and a specified evolution time.
@@
11 np.sqrt(p) * np.identity(2),
12 np.sqrt(1-p) * sz
@@
By design the ParamChannel class represents a discrete time quantum evolution. Thus to create its objects from $\mathcal{L}_{\omega},\dot{\mathcal{L}}_{\omega}$ one needs to move from continuous to discrete time regime by integrating over specified evolution time $t$. This can be done using choi_from_lindblad function from qtools which performs this integration for time-independent Lindbladian specified either as a function or a pair of Hamiltonian and a list of jump operators:
1# continuing previous example
2
3# Lindbladian and its derivative.
4omega = 0.0 # Parameter.
5t = 1 # Time.
6# Dephasing strength:
7gamma = -np.log(2*p-1)/t
8# Hilbert space dimension
9dim = 2
10
11def lindblad(rho):
@@
20
21choi, dchoi = choi_from_lindblad(
22 lindblad, dlindblad, t, dim=dim
23)
24channel5 = ParamChannel(
25 choi=choi, dchoi=dchoi
26)
27
28# or
29
30# Hamiltonian
31H = 0.5 * omega * sz
32# Jump operators rescaled by sqrt(gamma)
33Ls = [np.sqrt(gamma/2) * sz]
34# Derivative of the Hamiltonian
35dH = 0.5 * sz
36# Derivative of the jump operators
37dLs = [np.zeros_like(L) for L in Ls]
38choi, dchoi = choi_from_lindblad(
39 (H, Ls), (dH, dLs), t
40)
41channel6 = ParamChannel(
42 choi=choi, dchoi=dchoi
43)
Objects of the ParamChannel class can be used to compute $\rho_{\theta}=\Lambda_{\theta}(\rho)$ and $\dot{\rho}_{\theta}=\frac{d}{d\theta}\Lambda_{\theta}(\rho)$, e.g.:
@@
or to obtain the CJ operator from Kraus operators and vice versa, e.g.:
@@
3.2 Single-channel QFI optimization
Figure 4: (A) Diagram of a strategy with a single parameter-dependent channel $\Lambda_{\theta}$ and ancilla $\mathcal{A}$. Values of QFI for (B) dephasing and (C) amplitude damping, for various values of $p$ and for different methods: MOP - black $\times$, ISS with $d_{\mathcal{A}}=1$ - blue dashed line, ISS with $d_{\mathcal{A}}=2$ - red dotted line.
After defining the parameter-dependent channel using the ParamChannel class, we may now easily compute the single-channel QFI (9), including the possible use of an ancilla, see Fig. 4(A). The channel QFI can be computed using one of the two functions:
  • iss_channel_qfi for the ISS method,
  • mop_channel_qfi for the MOP method.
The function iss_channel_qfi takes two arguments: the channel whose QFI is computed and the dimension of the ancillary system. It returns five items: the optimized QFI, a list of pre-QFI values per algorithm iteration, the density matrix of the optimal input state, the SLD matrix, and the information whether the optimization converged successfully.